# Forecasting intraday foreign exchange volatility with functional GARCH approaches

### v2

@@ 2 Functional GARCH-type models
The intraday return curves can be expressed as $y_{t}(u)=c(u)+\widetilde{y}_{t}(u)$, where $c(u)$ represents the functional mean, and the mean-zero curve $\widetilde{y}_{t}(u)$ is the focus of interest. Given the increased prevalence of volatility clustering in data like $\widetilde{y}_{t}(u)$, Hörmann et al. (2013) first propose and establish the stationary solution of the FARCH(1) model for conditional volatility of $\widetilde{y}_{t}(u)$. However, this model has been shown to inadequately account for the conditional heteroscedasticity observed in financial assets (Rice et al., 2020b). Aue et al. (2017) and Cerovecki et al. (2019) generalise the FARCH model and specify the conditional volatility of $\widetilde{y}_{t}(u)$ following the recursion of,
[…]
@@
where $\sigma^{2}_{t}(u)$ is a latent conditional volatility function. The error term $\epsilon_{t}(u)$ is assumed to be a weakly dependent stationary process with $\text{E}(\epsilon_{t}(u))=0$ and $\text{E}(\epsilon^{2}_{t}(u))=1$, and $\text{E}[\epsilon^{2}_{0}(u)]=1$ for all $\epsilon\in[0,1]$. It is needed for the identifiability condition. The coefficient function $\omega(u)\in\mathcal{L}^{2}[0,1]_{*}^{+}$, and the kernels of the coefficient operators, $\alpha_{j}(u,v)$ and $\beta_{k}(u,v)$ are elements in $\mathcal{L}^{2}[0,1]^{+}\times\mathcal{L}^{2}[0,1]^{+}$, for $1\leq j\leq q$ and $1\leq k\leq p$. We refer interested readers to Cerovecki et al. (2019) for the stationary solution and the consistent estimation method of the FGARCH$(p,q)$ model. Suitable lag orders $p$ and $q$ can be found using the goodness-of-fit test for the conditional heteroscedasticity (Rice et al., 2020b).
@@
Another parsimonious framework that models conditional volatility dependence is the FGARCH-X model. Given empirical evidence, Rice et al. (2020b) suggest that the FGARCH(1,1) model is generally adequate to explain the short-range conditional heteroscedasticity in most data derived from financial markets. Rice et al. (2023) extends the FGARCH(1,1) to an FGARCH-X model that uses an exogenous covariate X to add explanatory ability to the rich dynamic of conditional heteroscedasticity. The model specifies the conditional volatility $\sigma_{t}^{2}(u)$:
[…]
@@
where $[\omega(u)+\boldsymbol{\gamma}(x_{t-1})(u)]\in\mathcal{L}^{2}[0,1]^{+}_{*}$ for all $t$, and the kernel of the coefficient operators $\alpha(t,s)$, $\beta(t,s)$ and $\gamma(t,s)\in\mathcal{L}^{2}[0,1]^{+}\times\mathcal{L}^{2}[0,1]^{+}$. The model has a stationary solution if the exogenous covariate is weakly stationary. According to Rice et al. (2023), the long-range conditional heteroscedasticity may be accounted for by incorporating a long-range dependent covariate X. Models (2.1) and (2.2) serve as forecasting models once a dataset is fitted and the model coefficients are estimated.
@@
To consistently estimate FGARCH-type models, two methods have been documented in the literature, namely ordinary least squares (OLS) (Aue et al., 2017) and quasi-maximum likelihood estimation (QMLE) (Cerovecki et al., 2019). Given the estimation efficiency advantages outlined in Cerovecki et al. (2019), we choose to estimate our models using the QMLE approach. Like all functional time-series models, the FGARCH-type models are intrinsically evaluated in an infinite-dimensional Hilbert space, meaning that they can only be estimated in a finite-dimensional space by way of dimension reduction. Following Cerovecki et al. (2019), we project the infinite-dimensional conditional volatilities $\sigma_{t}^{2}(u)$ into a finite $K$-dimensional subspace of $\mathcal{L}^{2}[0,1]^{+}$. Let $\{\psi_{1},\psi_{2},\dots,\psi_{K}\}\in\mathcal{L}^{2}[0,1]^{+}$, in which we denote $K$ as known linearly independent basis functions. We then project the kernel operators in models (2.1) and (2.2) onto these bases. For example, for the model (2.2) we have,
@@
Correspondingly, we will need to estimate the parameters $\theta=\text{vec}(\boldsymbol{D},\boldsymbol{A}_{1},\boldsymbol{B}_{1})$ for the models (2.1) when $p=q=1$. The QMLE estimator in Cerovecki et al. (2019) solves the optimisation
@@
The conditional variance $\widetilde{\sigma}_{t}^{2}(u)$ for model (2.1) can be deduced similarly. We rely on the regularity assumptions of Cerovecki et al. (2019, Theorem 2) to guarantee that QMLE is a strongly consistent estimator of $\widehat{\theta}_{N}$.
Importantly, to make the estimation feasible, we must decide on an optimal number for $K$ and the basis functions $\psi_{l}$. Existing methods include the use of goodness-of-fit tests (Shibata, 1981), cross-validation (Bathia et al., 2010), total variation explanation (Cattell, 1966), and the eigenvalue ratio method (Lam and Yao, 2012). The first two methods are based on a grid search principle and can be computationally burdensome. The total variation explanation method determines $K$ by choosing an adequate number of basis functions to cumulatively explain a certain proportion of the variation, for example, 90% of the total variations of our curve data. Lastly, the eigenvalue ratio method (Lam and Yao, 2012; Li et al., 2020) selects $K=l$ via
@@
where $\overline{L}$ is a prespecified positive integer and $\widehat{\lambda}_{l}$ is the $l$th largest eigenvalue associated with eigenfunction $\widehat{\phi}_{l}$. In this study, we use the eigenvalue ratio method. In choosing a suitable $\psi_{l}$, various types of data-driven methods are discussed in the literature. Cerovecki et al. (2019) suggest using the truncated FPCA, while Rice et al. (2023) propose functional sparse and non-negative basis and truncated predictive factor methods to decompose $\{\widetilde{y}_{t}^{2}(u)\}$.
@@
Note that, in addition to the FGARCH-type models, functional time series forecasting can also be performed using the FPCA approach of Aue et al. (2015), where the dynamics of functional scores are fitted and predicted by univariate or multivariate time series models. Simple adaptations for forecasting the conditional volatility of intraday return curves do not show clear empirical superiority in our unreported results. Therefore, we focus on FGARCH-type models, which benefit from the ability to capture the dynamics of the entire intraday return curves.
3 FX data and preliminary analysis
This section conducts preliminary analyses to understand the properties of our dataset and then shows the limitations of existing FGARCH-type models in forecasting FX intraday volatility. We consider the three most traded currencies against USD, namely EUR, GBP, and JPY. These currencies constituted over 75% of the turnover of OTC FX instruments in 2019, according to the Bank for International Settlements. [^2] Bank of International Settlements, https://stats.bis.org/statx/srs/table/d11.3 For each currency, the intraday spot, bid, and ask rates are collected at a 5-minute frequency from a collection of databases, including Morningstar, FXCM, TenFore and Thomson Reuters. The sample ranges from 07-January-2014 to 30-September-2020. Given the OTC nature of FX markets, our sample covers U.S. trading hours from midnight on Monday until the close on Thursday night. Intraday data on Sunday and Friday are truncated from our sample as they do not form 24-hour trading sessions. Consequently, each FX rate forms an information set of 1,275 trading days, each with 288 intraday grid points. In an unreported analysis, we also include forecasting during the U.S. main trading session from 7:00 to 17:00, Monday to Friday, spanning 1,722 trading days with 108 daily grid points each day. The results show consistency with the dataset used in this study.
$y_{t}(u)=[\ln P_{t}(u)-\ln P_{t-1}(1)]\times 100,$ (3.1)
where $P_{t}(u)$ denotes the spot rate of either currency on day $t$ evaluated at intraday time $u$, $u\in[0,1]$, and $P_{t-1}(1)$ is the closing spot rate on day $t-1$. Functional data objects can be obtained using smoothing techniques (Ramsay and Silverman, 2006, Chapter 3) or simple linear interpolation under the condition of continuity (Rice et al., 2020a). The OCIDR curves provide a continuous trajectory, so they are suitable for casting as functional time series objects while accounting for the overnight effect (Hansen and Lunde, 2006b). Desirably, forecasting the volatility of OCIDR curves includes inter-daily volatility prediction because OCIDR curves end at daily closing returns, defined as $y_{t}=\ln P_{t}(1)-\ln P_{t-1}(1)$.
Additionally, we obtain intraday bid-ask spread curves. The inclusion of the bid-ask spread is motivated by Evans and Rime (2019), who point out that, although the FX rate is commonly represented as the midpoint between the best bid and ask price, in reality, all FX trades are executed at either the bid or ask price. This bid or ask trading dynamic also brings about imperfect measurement in determining the variance of return, further motivating our functional characterisation of the intraday data. The use of bid-ask spreads has been widely discussed in the volatility forecasting literature. For example, Bollerslev and Melvin (1994) state that bid-ask spreads act as a measure of liquidity and have a significant relationship with FX rate volatility. Incorporating bid-ask spread data also helps address the issue of bid-ask bounce, which might be misinterpreted as volatility arising from price movements. As a measure of liquidity, we calculate the intraday bid-ask spread curves (IBAS) through
$\text{IBAS}_{t}(u)=\text{ask}_{t}(u)-\text{bid}_{t}(u),\quad u\in[0,1].$
To achieve stationarity and generate an information flow that is compatible with the OCIDR curve, we calculate the overnight cumulative IBAS (OCIBAS) curve using $\text{OCIBAS}_{t}(u)=\text{IBAS}_{t}(u)-\text{IBAS}_{t-1}(1)$. The OCIBAS curve is also assumed to be square integrable with sample path in $\mathcal{L}^{2}[0,1]$. Intuitively, higher values of the $\text{OCIBAS}_{t}(u)$ curve represent lower liquidity, and vice versa.
Figure 3.1 displays the plots of the OCIDR and OCIBAS curves for USD/EUR. Each OCIDR curve records the evolution of return movements over the trading day. The plot of the OCIBAS curves shows that they spike during opening and closing trading hours, implying less liquidity during these trading sessions. To understand the properties of these curve processes, we deploy the KPSS-type stationary test (Horváth et al., 2014), the static FPCA-type normality test (Górecki et al., 2018), the independence test (Kokoszka et al., 2017), and the conditional heteroscedasticity test (Rice et al., 2020b). Table 3.1 documents the p-values for each statistical test on the first training sample, which runs from 03-January-2014 to 31-December-2016. We limit our testing to the first training sample for in-sample analysis throughout the paper, and to ensure the robustness of our reported results, we usually replicate the analysis for other training samples. Table 3.1 indicates that the OCIDR curves are weakly stationary, non-Gaussian, serially uncorrelated, and conditionally heteroscedastic functional time series. Interestingly, the properties of the OCIDR curves uncovered here for the FX market are consistent with intraday return curves in equity and commodity markets (see, e.g., Rice et al., 2023).
Table 3.1: $P$-values of statistical tests on Overnight Cumulative Intraday Return (OCIDR) curves and Overight Cumulative Intraday Bid-Ask Spread (OCIBAS) curves, as defined in Section 3, for our three currency pairs, USD/EUR, USD/GBP, and USD/JPY. The results are calculated for 07-January-2014 to 17-May-2017, the first training sample period of our data. We test for stationarity, normality, autocorrelation, and heteroscedasticity with the $\widehat{a}_{x}$ and $\widehat{a}_{x^{2}}$/$\widehat{a}_{|x|}$ representing the short- and long-range dependency memory parameter estimates obtained using the Local Whittle (LW) estimator. Lags of 1, 5, 10, and 20 days are specified for the autocorrelation and heteroscedasticity tests.
Stationarity Normality Autocorrelation Heteroscedasticity $\widehat{a}_{x}$ $\widehat{a}_{x^{2}}$ Lag 1 5 10 20 1 5 10 20 $\mbox{OCIDR}_{\text{USD/EUR}}$ 0.22 0.00 0.71 0.61 0.43 0.60 0.04 0.00 0.00 0.00 0.02 0.20 $\mbox{OCIDR}_{\text{USD/GBP}}$ 0.60 0.00 0.64 0.68 0.56 0.40 0.06 0.00 0.00 0.00 -0.04 0.33 $\mbox{OCIDR}_{\text{USD/JPY}}$ 0.68 0.00 0.19 0.10 0.28 0.25 0.00 0.02 0.04 0.02 0.06 0.13 $\widehat{a}_{|x|}$ $\mbox{OCIBAS}_{\text{USD/EUR}}$ 0.11 0.00 0.37 0.25 0.36 0.37 0.96 0.09 0.01 0.00 0.11 0.12 $\mbox{OCIBAS}_{\text{USD/GBP}}$ 0.04 0.00 0.14 0.37 0.26 0.20 0.95 0.07 0.00 0.00 0.18 0.19 $\mbox{OCIBAS}_{\text{USD/JPY}}$ 0.07 0.00 0.61 0.83 0.34 0.34 0.00 0.00 0.00 0.00 0.10 0.27
We also investigate the long-range dependence structure of the level and squares of the FX curve sequences. According to Li et al. (2020, 2021), the memory parameter $a$ of stationary curves can be estimated either via a semiparametric R/S method or a feasible local Whittle (LW) estimator. Taking into account the efficiency of the estimator, we apply the latter method as suggested by Li et al. (2021), with the LW estimator $\widehat{a}_{x}$ denoting the memory parameter of the level curves, and $\widehat{a}_{x^{2}}$ denoting the memory parameter of the squared process. The estimates of $\widehat{a}_{x}$ and $\widehat{a}_{x^{2}}$ show that the levels of the OCIDR curves are short-range dependent, while the squared processes are long-range dependent. The absolute processes $\widehat{a}_{|x|}$ of OCIBAS curves are also long-range dependent.
Furthermore, it has been documented that co-movements between the excess returns of a variety of internationally traded financial assets are widespread (see, e.g. Bekaert, 1995). These studies motivate a series of work on forecasting daily volatility in the FX market. Recently, theoretical bases for understanding cross-currency co-movements have been built by Gabaix and Maggiori (2015) and Colacito et al. (2018), with Gabaix and Maggiori (2015) further highlighting the role of co-movements in modelling FX volatility. This cross-dependence structure between related currency pairs may also add predictability to the return volatility in FX markets. Therefore, we seek to understand the interdependency between currency pairs by estimating the sample covariance operator between their OCIDR curves. Figure 3.2 displays the estimated covariance operator between the various pairwise OCIDR curves.
The results show that all three pairs of OCIDR curves are correlated, with a particularly high correlation between USD/EUR and USD/GBP. The source of this cross-dependency can be attributed to shocks to the U.S. economy and U.S. interest rates, which, given the base currency of USD, impact all FX rates we consider. Eichenbaum and Evans (1995) provide empirical evidence that U.S. monetary policy shocks, such as interest rate changes, have significant spillover effects on exchange rates globally, highlighting how U.S. economic shifts create dependencies among various currencies. Hausman and Wongswan (2011) examine the impact of U.S. Federal Reserve announcements, particularly those concerning interest rates, on global asset prices, including exchange rates. This study supports the idea that cross-dependencies in currency markets often arise from U.S. economic shocks. Thus, the empirical finding of strong covariance motivates our position that accounting for cross dependence may enhance volatility modelling and forecasting.
4 Basis function specification
@@
Based on the nature of the FX intraday return curves described in Section 3, it is clear that the models presented in Section 2 struggle to capture the stylised features to enhance predictions of the conditional volatility of OCIDR curves. In this section, we show that this useful information can still be used by selecting more nuanced basis functions $\psi_{l}$ in  (2.3). As a crucial component of its dimension reduction procedure, basis selection drives the estimation and forecasting performance of FGARCH-type models. Usually, one can apply static (Cerovecki et al., 2019; Paparoditis and Shang, 2023) or dynamic FPCA methods (Bathia et al., 2010; Hörmann et al., 2015; Shang and Kearney, 2022) to obtain basis functions. However, these methods overlook the rich cross-dependence of $\{y_{t}^{2}(u)\}$.
@@
This section introduces two new data-driven approaches to decompose basis functions $\psi_{l}$ from the square OCIDR $\{y_{t}^{2}(u)\}$ curves. We decompose $\{y_{t}^{2}(u)\}$ here in the hope that the functional scores $y^{2}_{t}(u),\psi_{l}(u)\right>$ preserve intraday conditional volatility dependence after we project the conditional volatility of $y_{t}(u)$ onto these basis functions. Note that intraday seasonal patterns are also absorbed into the estimated functional principal components and their associated scores. Therefore, modelling the dynamics of the scores considers intraday seasonality.
4.1 Multi-level FPCA

### v3

@@ 3.1 Functional GARCH-type models
The intraday return curves can be expressed as $y_{t}(u)=c(u)+\widetilde{y}_{t}(u)$, where $c(u)$ represents the functional mean, and the mean-zero curve $\widetilde{y}_{t}(u)$ is the focus of interest. Given the increased prevalence of volatility clustering in data like $\widetilde{y}_{t}(u)$, Hörmann et al., (2013) first propose and establish the stationary solution of the FARCH(1) model for the conditional volatility of $\widetilde{y}_{t}(u)$. However, this model has been shown to inadequately account for the conditional heteroskedasticity observed in financial assets (Rice et al., 2020b, ). Aue et al., (2017) and Cerovecki et al., (2019) generalise the FARCH model and specify the conditional volatility of $\widetilde{y}_{t}(u)$ following the recursion of,
$\begin{split}&\widetilde{y}_{t}(u)=\sigma_{t}(u)\varepsilon_{t}(u),\quad u\in[0,1],\quad\mbox{ }1\leq t\leq N,\\ &\sigma^{2}_{t}(u)=\omega(u)+\sum_{j=1}^{q}\boldsymbol{\alpha}_{j}(\widetilde{y}^{2}_{t-j})(u)+\sum_{k=1}^{p}\boldsymbol{\beta}_{k}(\sigma^{2}_{t-k})(u),\\ &\qquad\mbox{ }=\omega(u)+\sum_{j=1}^{q}\int\alpha_{j}(u,v)\widetilde{y}_{t-j}^{2}(v)dv+\sum_{k=1}^{p}\int\beta_{k}(u,v)\sigma_{t-k}^{2}(v)dv,\end{split}$ (3.1)
@@
where $\sigma^{2}_{t}(u)$ is a latent conditional volatility function. The error term $\epsilon_{t}(u)$ is assumed to be a stationary process with weak dependence with $\text{E}(\epsilon_{t}(u))=0$ and $\text{E}(\epsilon^{2}_{t}(u))=1$, and $\text{E}[\epsilon^{2}_{0}(u)]=1$ for all $\epsilon\in[0,1]$. It is needed for the identifiability condition. The coefficient function $\omega(u)\in\mathcal{L}^{2}[0,1]_{*}^{+}$, and the kernels of the coefficient operators, $\alpha_{j}(u,v)$ and $\beta_{k}(u,v)$ are elements in $\mathcal{L}^{2}[0,1]^{+}\times\mathcal{L}^{2}[0,1]^{+}$, for $1\leq j\leq q$ and $1\leq k\leq p$. We refer interested readers to Cerovecki et al., (2019) for the stationary solution and the consistent estimation method of the FGARCH(p,q)(p,q) model. Appropriate lag orders $p$ and $q$ can be selected using a goodness-of-fit test for conditional heteroskedasticity (Rice et al., 2020b, ), which shows that the FGARCH(1,1) model adequately captures the conditional heteroskedasticity of intraday return curves for many financial assets.
@@
Another parsimonious framework that models conditional volatility dependence is the FGARCH-X model. Rice et al., (2023) extends the FGARCH(1,1) to an FGARCH-X model that uses an exogenous covariate X to add explanatory ability to the rich dynamic of conditional heteroskedasticity. The model specifies the conditional volatility $\sigma_{t}^{2}(u)$:
$\begin{split}&\sigma_{t}^{2}(u)=\omega(u)+\boldsymbol{\alpha}(\widetilde{y}^{2}_{t-1})(u)+\boldsymbol{\beta}(\sigma^{2}_{t-1})(u)+\boldsymbol{\gamma}(x_{t-1})(u),\\ &\qquad\mbox{ }=\omega(u)+\int\alpha(u,v)\widetilde{y}_{t-1}^{2}(v)dv+\int\beta(u,v)\sigma_{t-1}^{2}(v)dv+\int\gamma(u,v)x_{t-1}(v)dv,\end{split}$ (3.2)
@@
where $\omega(u)+\boldsymbol{\gamma}(x_{t-1})(u)]\in\mathcal{L}^{2}[0,1]^{+}_{*}$ for all $t$, and the kernel of the coefficient operators $\alpha(t,s)$, $\beta(t,s)$ and $\gamma(t,s)\in\mathcal{L}^{2}[0,1]^{+}\times\mathcal{L}^{2}[0,1]^{+}$. The model has a stationary solution if the exogenous covariate is weakly stationary. According to Rice et al., (2023), the long-range conditional heteroskedasticity may be accounted for by incorporating a long-range dependent covariate X. Models (3.1) and (3.2) serve as forecasting models once a dataset is fitted and the model coefficients are estimated.
@@
To consistently estimate FGARCH-type models, two methods have been documented in the literature, namely ordinary least squares (OLS) (Aue et al.,, 2017) and quasi-maximum likelihood estimation (QMLE) (Cerovecki et al.,, 2019). Given the advantages of efficiency estimation outlined in Cerovecki et al., (2019), we choose to estimate our models using the QMLE approach. Like all functional time-series models, the FGARCH-type models are intrinsically evaluated in an infinite-dimensional Hilbert space, meaning that they can only be estimated in a finite-dimensional space by way of dimension reduction. Following Cerovecki et al., (2019), we project the infinite-dimensional conditional volatilities $\sigma_{t}^{2}(u)$ into a finite $K$-dimensional subspace of $\mathcal{L}^{2}[0,1]^{+}$. Let $\{\psi_{1},\psi_{2},\dots,\psi_{K}\}\in\mathcal{L}^{2}[0,1]^{+}$, in which we denote $K$ as known linearly independent basis functions. We then project the kernel operators in models (3.1) and (3.2) onto these bases. For example, for the model (3.2) we have,
@@
Consequently, we will need to estimate the parameters $\theta=\text{vec}(\boldsymbol{D},\boldsymbol{A}_{1},\boldsymbol{B}_{1})$ for the models (3.1) when $p=q=1$. The QMLE estimator in Cerovecki et al., (2019) solves the optimisation
@@
The conditional variance $\widetilde{\sigma}_{t}^{2}(u)$ for model (3.1) can be deduced similarly. We rely on the regularity assumptions of Cerovecki et al., (2019, Theorem 2) to guarantee that QMLE is a strongly consistent estimator of $\widehat{\theta}_{N}$.
Importantly, to make estimation feasible, we must decide on an optimal number for $K$ and the basis functions $\psi_{l}$. Existing methods include the use of goodness-of-fit tests (Shibata,, 1981), cross-validation (Bathia et al.,, 2010), total variation explanation (Cattell,, 1966), and the eigenvalue ratio method (Lam and Yao,, 2012). The first two methods are based on a grid search principle and can be computationally burdensome. The total variation explanation method determines $K$ by choosing an adequate number of basis functions to cumulatively explain a certain proportion of the variation, for example, 90% of the total variations of our curve data. Lastly, the eigenvalue ratio method (Lam and Yao,, 2012; Li et al.,, 2020) selects $K=l$ via
@@
where $\overline{L}$ is a prespecified positive integer and $\widehat{\lambda}_{l}$ is the $l$th largest eigenvalue associated with the eigenfunction $\widehat{\phi}_{l}$. In this study, we use the eigenvalue ratio method. In choosing a suitable $\psi_{l}$, various types of data-driven methods are discussed in the literature. Cerovecki et al., (2019) suggest using the truncated FPCA, while Rice et al., (2023) propose functional sparse and non-negative basis and truncated predictive factor methods to decompose $\{\widetilde{y}_{t}^{2}(u)\}$.
@@
Note that in addition to FGARCH-type models, functional time series forecasting can also be performed using the FPCA approach of Aue et al., (2015), where the dynamics of functional scores are fitted and predicted by univariate or multivariate time series models. Simple adaptations for forecasting the conditional volatility of intraday return curves do not show clear empirical superiority in our unreported results. Therefore, we focus on FGARCH-type models, which benefit from the ability to capture the dynamics of the entire intraday return curves.
3.2 Basis function specification
@@
Based on the nature of the FX intraday return curves described in Section 2, it is clear that the models presented in Section 3.1 struggle to capture stylised features to improve predictions of the conditional volatility of the OCIDR curves. In this section, we show that this useful information can still be used by selecting more nuanced basis functions $\psi_{l}$ in (3.3). As a crucial component of its dimension reduction procedure, basis selection drives the estimation and forecasting performance of FGARCH-type models. Usually, one can apply static (Cerovecki et al.,, 2019; Paparoditis and Shang,, 2023) or dynamic FPCA methods (Bathia et al.,, 2010; Hörmann et al.,, 2015; Shang and Kearney,, 2022) to obtain basis functions. However, these methods overlook the rich cross-dependence of $\{y_{t}^{2}(u)\}$.
@@
This section introduces two new data-driven approaches to decompose basis functions $\psi_{l}$ from the square OCIDR $\{y_{t}^{2}(u)\}$ curves. We decompose $\{y_{t}^{2}(u)\}$ here in the hope that the functional scores $\left<y^{2}_{t}(u),\psi_{l}(u)\right$> preserve intraday conditional volatility dependence after we project the conditional volatility of $y_{t}(u)$ onto these basis functions. Note that intraday seasonal patterns are also absorbed into the estimated functional principal components and their associated scores. Therefore, modelling the dynamics of the scores considers intraday seasonality.
3.2.1 Multi-level FPCA